OpenStudy (anonymous):

7 sin2x - 14 sin x + 2 = -5

OpenStudy (ajprincess):

Is it \(\sin^2x\)?

OpenStudy (ajprincess):

@needinghelpyo

OpenStudy (anonymous):

can you show me how you got that cause i dont get it at all haha

OpenStudy (ajprincess):

is ur question \(7\ sin^2x - 14\ sin x + 2 = -5\)?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

\[7{sin}^2x-14sinx+2+5=0\] \[7{sin}^2x-14sinx+7=0\] U can factor out 7. Then u will get \[7({sin}^2x-2sinx+1)=0\] When u divide by 7 both sides u will get \[{sin}^2x-2sinx+1=0\] Nw factorise this and solve for sinx. Can u do that @needinghelpyo?

OpenStudy (anonymous):

or let \(u=sinx\) might be easier with a sub :)

OpenStudy (ajprincess):

yup u r right @omniscience:)

OpenStudy (anonymous):

then would it be sin^2x=1/2

OpenStudy (ajprincess):

hmm no. As @omniscience said if u substitute u in place of sinx then u will get \(u^2-2u+1=0\) \(u^2-u-u+1=0\) \(u(u-1)-1(u-1)=0\) \((u-1)(u-1)=0\) \((u-1)^2=0\) \(u-1=0\) \(u=1\) \(sinx=1\) \(x=sin^{-1}1\)

OpenStudy (ajprincess):

Getting it @needinghelpyo?

OpenStudy (anonymous):

yes now i do thank you so much

OpenStudy (ajprincess):

yw:)